OpenStudy (anonymous):

Rewrite the given equation using the substitutions x=rcos(theta) & y=rsin(theta). Simplify. x^2 + y^2 + 3x = 0 I know exactly how to solve it but what do the r's mean? :/

OpenStudy (anonymous):

So how would I solve? I thought I could simplify rcos(theta)^2 + rsin(theta)^2 to 1. Can I still?

OpenStudy (loser66):

nope, you just replace x, y into the original one. and solve it. that's it

OpenStudy (akashdeepdeb):

Write it like this! \[r^2 (\sin \theta)^2 + r^2 (\cos \theta)^2 + 3r \sin \theta = 0\] Or \[r^2 + 3r \sin \theta = 0\] The r here is actually a way of representing a complex number in short! Have you learnt imaginary numbers? :)

OpenStudy (loser66):

the last term is 3x, not 3y friend

OpenStudy (akashdeepdeb):

My bad! :) Write it like this then @aodfuj2 r2+3rcosθ=0 Got it? :)

OpenStudy (anonymous):

so how did you get r2 from sin(theta)2 + cos(theta)2? Sorry I'm trying to understand for my next problems.

OpenStudy (akashdeepdeb):

x = rcosθ So x2 = (rcosθ)^2 = r^2 . cos^2 θ Got it? And then take r^2 common and use sin^2 θ + cos^2 θ = 1 Understood? :)

OpenStudy (anonymous):

Got it. Thank you!

OpenStudy (akashdeepdeb):

:)